# Conditional Dependence via U-Statistics Pruning

### v1

@@ IV Conditional dependence via U-Statistics
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)$ (14)
@@
where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (12)) is the cross-covariance matrix conditioned to a specific value $z$ of a confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution. Note that (14) is the most general definition of conditional cross-covariance that is valid for any statistics, while the Schur complement expression is only justified for Gaussian vectors [9, Ch. 4]. With the goal of deriving an estimator for (14), and following a similar rationale to $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$, we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$. Given that integrating all values of $\mathsf{z}$ is equivalent to doing so for $\mathring{\mathsf{z}}=0$, i.e. $\mathsf{z}_{1}=\mathsf{z}_{2}$, the expectation in (14) can be alternatively expressed as
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\int_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (15)
@@
where $\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and since $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal. In consequence, conditioning the covariance matrix with respect to $\mathsf{z}$ is equal to conditioning it with respect to $\mathring{\mathsf{z}}=0$. Therefore, the result in (IV) suggests letting the potential confounder data control the pruning of the incomplete U-Statistics in (13). But, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, the data control should be based on merely small values of $|\mathring{z}|$. Note that the procedure followed here is the same followed in [12].
@@
With the intention of choosing the data pairs for the pruning based on $|\mathring{z}|$, we define the samples of $\mathring{z}$ as follows:
$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),$ (16)
@@
where $z(l)$ and $z(l^{\prime})$ are i.i.d. samples drawn from $\mathsf{z}$ with $l\neq l^{\prime}$. Then, we let the sorting of $|\mathring{z}_{l,l^{\prime}}|$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in (11). Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$, i.e.
$K_{\alpha}\triangleq\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (17)
@@
being $1\leq\alpha\ll(L-1)$ a tuning hyper-parameter. While $\alpha=1$ ensures that only very small values of $|\mathring{z}_{l,l^{\prime}}|$ are considered, in the other extreme of $\alpha=L-1$ the U-Statistic becomes complete and there is no conditioning at all, thus encountering the HSIC as a particular case as in (7). A natural trade-off then emerges: while $\alpha$ close to $1$ is desirable for the conditioning idea (IV) to work properly, higher values of $\alpha$ will avoid excessive pruning and provide sufficient statistical accuracy in (13) with $K=K_{\alpha}$. Remarkably, the selection of the hyper-parameter $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as was shown in [12] under the correlation measure framework between a pair of vectors, and will be seen later on by a numerical example.
@@ IV-A Conditional HSIC
Now that we have determined the sorting and pruning of the data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (13):
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (18)
where the circularity of the trace has been used. To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrices $\mathring{\mathbf{U}}$ and $\mathring{\mathbf{V}}$ as follows:
$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\qquad\mathring{\mathbf{V}}=\tfrac{1}{\sqrt{2}}(\mathbf{V}_{1}-\mathbf{V}_{2}),$ (19)
where
$\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K_{\alpha})}],\;\;\mathbf{V}_{a}\triangleq[\mathbf{v}_{\mathrm{f}_{a}(1)},\dots,\mathbf{v}_{\mathrm{f}_{a}(K_{\alpha})}],$ (20)
for $a=\{1,2\}$. Accordingly, (18) is then rewritten as
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}$ (21)
$=\tfrac{1}{4K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
Taking the limit $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (21):
$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{x}_{a},\mathbf{x}_{a^{\prime}}}\right)\quad$ (22a)
$\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y}_{a},\mathbf{y}_{a^{\prime}}}\right),$ (22b)
where $\mathbf{S}\mathopen{}\mathclose{{}\left(\bm{\cdot},\bm{\cdot}}\right)$ is given in (9). Therefore, the new relevant data subsets for inferring conditional dependence are given by
$[\mathbf{x}_{a}]_{k}\triangleq x(f_{a}(k)),\quad[\mathbf{y}_{a}]_{k}\triangleq y(f_{a}(k)).$ (23)
Finally, the resulting C-HSIC can be expressed as follows:
@@
with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{2,1}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{2,1}$. Note that, as a result of the U-Statistics implementation, each entry of the new kernel-based matrices involves four data samples of the same source ($4$-tuples), in contrast to only the pairs typically involved in classical kernel methods. This fact, along with the lack of $\mathbf{P}$, are the main distinctive features of the C-HSIC (24) vs. HSIC (7).
V Numerical illustration
In order to test the proposed method, we aim at generating uncorrelated data with a controlled amount of co-information [17] (also called interaction information [18]), defined as
$\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})=\mathrm{I}(\mathsf{x};\mathsf{y})-\mathrm{I}(\mathsf{x};\mathsf{y}|\mathsf{z}),$ (25)
such that the potentials of the method can be clearly highlighted. Similar ideas on modeling conditional dependences can be found in [19]. Since co-information can be either positive or negative, two scenarios are studied: $\mathcal{M}^{+}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})>0$ and $\mathcal{M}^{-}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})<0$, whose variables are as follows:
@@ V Numerical illustration
The internal i.i.d. random variables are distributed as $\mathsf{a},\mathsf{b},\mathsf{c}\sim\mathcal{U}(0,\sqrt{3\text{}})$ (uniform), $\mathsf{v},\mathsf{w}\sim\mathcal{N}(0,1)$ (normal), and $\mathsf{p},\mathsf{q}\sim\mathrm{Bern}_{1/2}\{-1,1\}$ (equiprobable Bernoulli). Parameter $\gamma$ is the signal-to-noise ratio associated to the measurements and controls the total amount of absolute co-information. In model $\mathcal{M}^{+}$, $\mathsf{x}$ and $\mathsf{y}$ are dependent due to the influence of $\mathsf{a}$ in both, but they are conditionally independent, since knowing $\mathsf{z}$ (i.e. $\mathsf{a}$) implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$). By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in model $\mathcal{M}^{-}$, but they become conditionally dependent, since the knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$. In both models, $\mathsf{x}$, $\mathsf{y}$ and $\mathsf{z}$ are mutually uncorrelated due to the multiplicative effect of mutually independent variables $\mathsf{p}$ and $\mathsf{q}$, so correlation measures are unable to discover any data association.
Regarding the choice of kernel, the universal Gaussian kernel is used, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$. Note that the kernel length scales in proportion with the sample standard deviation $\hat{\sigma}$ of the data and in inverse proportion with the data size $L$. The reader is referred to [7, Appendix D][20, 21] for a justification of the typical power law $O(L^{-1/5})$.
@@
Figure 1: Model $\mathcal{M}^{+}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
Figure 2: Model $\mathcal{M}^{-}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
@@
In Fig. 1 and Fig. 2, the measures of dependence are shown as a function of $\gamma$ for both models with a moderate hyper-parameter $\alpha=4$ from (17). This yields $8\%$ of the total data pairs for $L=100$ and $1.3\%$ for $L=600$ (and even less for higher $L$). For model $\mathcal{M}^{+}$ (Fig. 1), the C-HSIC correctly depicts small conditional dependence, while the HSIC confirms that marginal dependence is high for moderate $\gamma$ values. Conversely, Fig. 2 exhibits the capability of the C-HSIC to discover significant conditional dependence for moderate $\gamma$ values for model $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by the HSIC at any value of $\gamma$.
To get further insights on the proposed ideas, Fig. 1-(b) and Fig. 2-(b) show the performance of the C-HSIC when the pruning of data pairs in the U-Statistics is random, which means that the pair $(l(k),l^{\prime}(k))$ is not controlled by the confounder. It is seen that the C-HSIC produces a measure that mimics the marginal dependence one, thus confirming the main property of unbiasedness related to the general U-Statistics theory, and also showing its consequent increase in variance. It is also important to note that the selected value of $\alpha$, and the small percentage of data that yields this choice, confirms that the increase in variance induced by the pruning becomes small for sufficiently large data sizes due to Remark 1, making $\alpha$ a noncritical free hyper-parameter.

### v2

@@ IV Conditional dependence via U-Statistics
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z),$ (10)
@@
where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (8)) is the cross-covariance matrix conditioned to a specific value $z$ of confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution. With the goal of deriving an estimator of (10), we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$. Integrating over all values of $\mathsf{z}$ is equivalent to doing so over the events in which $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ take the same value, i.e. $\mathring{\mathsf{z}}=0$. Therefore, the expectation in (10) can be alternatively expressed as
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\iint_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\iint_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (11)
@@
where $\iint_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal. In consequence, conditioning with respect to $\mathsf{z}$ is equal to conditioning with respect to $\mathring{\mathsf{z}}=0$. Therefore, (IV) suggests to let the pruning of the incomplete U-Statistics in (9) be handled by the potential confounder. However, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, data control should be based on merely small values of $\lvert\mathring{z}\rvert$ [13].
@@
With the intention of choosing data pairs to prune according to $\lvert\mathring{z}\rvert$, we define the samples of $\mathring{\mathsf{z}}$ as follows:
$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),\qquad l\neq l^{\prime},$ (12)
@@
where $z(l)$ and $z(l^{\prime})$ are i.i.d. samples drawn from $\mathsf{z}$. Then, we let the sorting of $\lvert\mathring{z}_{l,l^{\prime}}\rvert$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in Sec. III. Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$ with
$K=\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (13)
@@
being $1\leq\alpha\ll(L-1)$ a tuning hyper-parameter. While $\alpha=1$ ensures that only very small values of $\lvert\mathring{z}_{l,l^{\prime}}\rvert$ are considered, the U-Statistic becomes complete and there is no conditioning at all for $\alpha=L-1$, thus yielding the HSIC as a particular case as in (7). Then a natural trade-off emerges: low values of $\alpha$ are desirable to provide strong conditioning of data, but may also lead to excessive pruning with low statistical accuracy in (9). Remarkably, the selection of $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as shown in [13] under the correlation measure framework between a pair of vectors. This will be further discussed in Sec. V with a numerical example.
@@ IV-A Conditional HSIC
Now that we have determined the sorting and pruning of data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (9):
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (14)
where the circularity of the trace has been used. To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrix $\mathring{\mathbf{U}}$ as follows:
$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\quad\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K)}],$ (15)
for $a=\{1,2\}$. The same is done for $\mathring{\mathbf{V}}$. Accordingly, (14) is then rewritten as
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=$ (16)
$\quad\tfrac{1}{4K^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
Taking the limit of $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (16):
$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}},\qquad\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}},$ (17)
whose elements are
$[\mathbf{K}_{a,a^{\prime}}]_{k,k^{\prime}}$ $=\kappa\bigl{(}x(\mathrm{f}_{a}(k))-x(\mathrm{f}_{a^{\prime}}(k^{\prime}))\bigr{)},$ (18a)
$[\mathbf{Q}_{a,a^{\prime}}]_{k,k^{\prime}}$ $=\kappa\bigl{(}y(\mathrm{f}_{a}(k))-y(\mathrm{f}_{a^{\prime}}(k^{\prime}))\bigr{)},$ (18b)
where $\kappa(\bm{\cdot})$ is the kernel function as in (6). Finally, the resulting C-HSIC can be expressed as follows:
@@
with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{1,2}^{\mathrm{H}}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{1,2}^{\mathrm{H}}$. Note that, as a result of the U-Statistics implementation, each entry of the new kernel-based matrices involves four data samples of the same source ($4$-tuples), in contrast to only the pairs typically involved in classical kernel methods. This fact, along with the lack of $\mathbf{P}$, are the main distinctive features of the C-HSIC (19) vs. HSIC (7). In contrast to other conditional dependence measures, (19) does not require any matrix inversion. Since $K$ is set to grow linearly with $L$ in (13), the computational complexity is $O(L^{2})$.
V Numerical illustrations
To test the proposed method, we aim at generating uncorrelated data with a controlled amount of conditional dependence. Two scenarios are studied, $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$, defined as follows:
@@ V Numerical illustrations
The internal i.i.d. random variables are distributed as $\mathsf{a},\mathsf{b},\mathsf{c}\sim\mathcal{U}(0,\sqrt{3\text{}})$ (uniform), $\mathsf{v},\mathsf{w}\sim\mathcal{N}(0,1)$ (normal), and $\mathsf{p},\mathsf{q}\sim\mathrm{Bern}_{1/2}\{-1,1\}$ (equiprobable Bernoulli). In $\mathcal{M}^{+}$, the pair $\{\mathsf{x},\mathsf{y}\}$ are dependent, i.e. their mutual information is greater than zero, but they are conditionally independent, since knowing $\mathsf{z}$ implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$). By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in $\mathcal{M}^{-}$, but become conditionally dependent, since knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$. Parameter $\gamma$ is the signal-to-noise ratio associated with the measurements and controls the total amount of unconditional dependence in $\mathcal{M}^{+}$ and conditional dependence in $\mathcal{M}^{-}$. In both models, $\mathsf{x}$, $\mathsf{y}$ and $\mathsf{z}$ are mutually uncorrelated due to the multiplicative effect of mutually independent variables $\mathsf{p}$ and $\mathsf{q}$, so correlation measures are unable to discover any data association. Similar ideas on modeling conditional dependencies can be found in [21], which are inspired by co-information [22, 23].
The universal Gaussian kernel is used as the kernel function of choice, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$ being $\hat{\sigma}$ the sample standard deviation. We use this expression because of its association with kernel density estimation, known to be related to kernel methods when estimating certain dependence measures [24, Ch. 2]. This connection also justifies the adoption of the power law $O(L^{-1/5})$ [25, Ch. 3][15, Appx. D].
(a) Model $\mathcal{M}^{+}$.
(b) Model $\mathcal{M}^{-}$.
@@
Figure 1: Dependence measures as a function of $\gamma$ (with $L=100$) and $L$ (with $\gamma=10$dB). C-HSIC measures conditional dependence while HSIC measures unconditional dependence. Markers denote the empirical average and bands indicate the standard deviation.
@@
Fig. 1 shows measures of dependence for both models with two $\alpha$ values from (13): a moderate $\alpha=4$ and a sixteenfold increase $\alpha=64$. This yields $4\%$ and $64.5\%$, respectively, of the total data pairs for $L=100$, or $0.7\%$ and $10.7\%$ for $L=600$ (and even less for higher $L$). For $\mathcal{M}^{+}$ (Fig. 1-(a)), C-HSIC correctly depicts conditional independence for an increasing value of $\gamma$, while HSIC confirms that marginal dependence is high as $\gamma$ increases. Conversely, Fig. 1-(b) exhibits the capability of C-HSIC to discover conditional dependence for moderate and high $\gamma$ values in $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by HSIC at any value of $\gamma$.
To get further insights on the proposed ideas, Fig. 1-(a.1) and Fig. 1-(b.1) also show the performance of C-HSIC when pruning data pairs is performed randomly in the U-statistic, i.e. the pair given by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ is not controlled by the confounder. As a result of not being properly conditioned, C-HSIC produces a measure of marginal dependence similar to HSIC, but with increased variance due to the pruning itself. It is also worth noting that random and orderly pruning perform similarly for low values of $\gamma$, corresponding to the regime where the confounder has no influence on the pair $\{\mathsf{x},\mathsf{y}\}$, and start to deviate as $\gamma$ increases. Additionally, it can be seen that the expected value deteriorates for $\alpha=64$, increasing in $\mathcal{M}^{+}$ and decreasing in $\mathcal{M}^{-}$. This is a consequence of a mild pruning for $L=100$ (where $\alpha=64$ is intentionally chosen to illustrate this effect), worsening the conditioning process and starting to behave as an unconditioned measure.
Finally, Fig. 1-(a.2) and Fig. 1-(b.2) illustrate the behavior of C-HSIC for multiple values of $L$ at $\gamma=10$dB. Model $\mathcal{M}^{+}$, which is conditionally independent, shows that C-HSIC correctly approaches zero as $L$ increases. Conversely, it tends to some nonzero value in $\mathcal{M}^{-}$. As mentioned above, $\alpha=64$ is using too many data pairs for small $L$ values and provides a degraded empirical average. In both models, $\alpha=4$ shows an increased variance with respect to $\alpha=64$ due to a more thorough pruning. However, this is reduced for sufficiently large data sizes due to Remark 1. Similarly, the gap between expected values for $\alpha=4$ and $\alpha=64$ is reduced as $L$ increases, albeit the former has the advantage of a lower computational complexity. These results show that $\alpha$ is a noncritical hyper-parameter, provided it is reasonably small.